OpenStudy (experimentx):

Evaluate:- \[ \lim_{n\to\infty} \left ( \sin \left( {\frac{n}{n^2+1^2}} \right ) +\sin \left ( {\frac{n}{n^2+2^2}} \right ) +\cdots+\sin \left ( {\frac{n}{n^2+n^2}} \right ) \right ) \] My guess is this converges to \( \pi \over 4 \)

OpenStudy (anonymous):

\[\lim_{n\to\infty}\sin\left(\frac{n}{n^2+1}\right)=0\\ \vdots\\\lim_{n\to\infty}\sin\left(\frac{n}{n^2+n^2}\right)=\lim_{n\to\infty}\sin\left(\frac{1}{2n}\right)=0\] I think it's 0.

OpenStudy (anonymous):

we write this as a sum \[\huge \lim_{n \rightarrow 0} \huge \sum_{k=1}^{n}\sin(\frac{ n }{ n^2+k^2 })\] \[\huge \lim_{n \rightarrow 0} \huge \sum_{k=1}^{n}\frac{1}{n}\sin \left(\frac{ 1 }{ 1+(\frac{ k }{ n })^2 }\right)\] evaluating the sum using integrals we have \[\Delta x=\frac{ b-a }{ n }=\frac{ 1 }{ n }\] \[x=k/n\] \[\huge \int\limits_{0}^{1} \sin \left(\ \frac{ 1 }{ 1+x^2 }\right)dx\]

OpenStudy (anonymous):

i am still trying to figure out the result

OpenStudy (anonymous):

oh i see the problem 1/n shud be in the parenthesis

OpenStudy (anonymous):

lets look at the question if there s no sin we have \[\int\limits \frac{ 1 }{ 1+x^2 }=\tan^{-1}x\] so between 0 and 1 its \[\pi/4\]

OpenStudy (experimentx):

yeah that's correct ...

OpenStudy (experimentx):

you should use the fact that sin(x) ~ x when x->0 and when n->oo, you get that thing.

OpenStudy (anonymous):

so we use the advantage of the continuity of the sign function and do the sum inside the sin

OpenStudy (experimentx):

yes ... that's what i thought ..